40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 50     Entry time: Fri Feb 5 05:13:09 2010
Author: Peter, Frank 
Type: Laser 
Category: RefCav 
Subject: everything back online 

we spent the whole night to re-align everything. By now everything is back online, both cavities were locked.for a short time. We still have the pointing problem. It's different now, seems to be smaller but is now almost the same for x and y. Will do this later today ....

The 10kHz peak in the old spectrum seems to be related to the crossover frequency of the fast actuator and the pc of the fss loop.

The PDH box is modified, all filter stages now have socket adapter boards to easily change the filter frequencies by changing the capacitor value within seconds to optimize everything. A zero for compensating the cavity pole is also installed.  A modulation input too.

ELOG V3.1.3-